# A Comprehensive Survey on Automatic Text Summarization with Exploration of LLM-Based Methods

### v2

@@ Abstract
The exponential growth of textual content on the internet, alongside vast archives of news articles, scientific papers, legal documents, and other domains, has made Automatic Text Summarization (ATS) increasingly important. ATS aims to create concise and accurate summaries, significantly reducing the effort required to process large volumes of text. Originating in the 1950s, ATS has evolved through several technical shifts, moving from statistical models to machine learning and deep learning approaches, and more recently, to pre-trained models. Previous surveys have focused on conventional ATS methods, which are often constrained by predefined generative paradigms. However, the advent of Large Language Models (LLMs) has introduced a paradigm-flexible approach to summarization. With their superior generative capabilities, in-context learning, and few-shot learning abilities, LLMs have demonstrated remarkable improvements in coherence, fluency, and overall summarization quality. In this survey, we provide a comprehensive review of both conventional ATS approaches and the latest advancements in LLM-based methods. Additionally, we propose a novel retrieval algorithm designed to efficiently collect relevant papers, which could be adapted for use in other types of surveys. Our contributions are threefold: (1) offering an up-to-date survey of ATS, (2) reviewing the latest LLM-based summarization methods, and (3) introducing a new retrieval algorithm for paper collection.
@@ 1 Introduction
The exponential growth of the World Wide Web has led to an overwhelming surge in textual data across various domains, including news articles, websites, user reviews, blogs, and social media. Additionally, vast text archives are available in specialized fields such as books, scholarly papers, legal documents, and biomedical records. This rapid expansion of data has far outpaced the ability of individuals to search, read, and process all relevant information. To address this challenge, the field of Automatic Text Summarization (ATS) has emerged, leveraging methods from multiple research areas such as Natural Language Processing (NLP) and Information Retrieval (IR), alongside advanced techniques like Machine Learning (ML), Deep Learning (DL), and Large Language Models (LLMs). ATS enables users to efficiently grasp the key ideas within texts, significantly reducing the time and effort required for document reading and comprehension.
@@
Historically, ATS has undergone several distinct eras of development. In the pre-2000s, early approaches were primarily statistical, leveraging surface-level features such as word frequency and sentence position to generate summaries [1, 2, 3]. From the 2000s to 2010s, machine learning and shallow neural network-based methods emerged, automating the feature extraction process but still heavily relying on domain-specific, handcrafted features [4, 5, 6, 7, 8]. Along with the introduction of deep learning in the 2010s, particularly with encoder-decoder architectures like RNNs [9] and LSTMs [10], enabled more nuanced, context-aware summarization by modeling sequences of words and capturing richer semantic information. During this period, two main paradigms dominated: extractive summarization, which focused on identifying key sentences or phrases from the original text [11, 12], and abstractive summarization, which generated new sentences to paraphrase the main ideas [13, 14]. However, these approaches were constrained by their inflexible generative paradigms, which were determined by the specific model architecture and training data. Moreover, abstractive summarization often faced challenges in maintaining content coherence and ensuring factual accuracy.
@@
The advent of Large Language Models (LLMs), such as GPT-3 and ChatGPT, represents a significant breakthrough in ATS by offering a paradigm-flexible approach. Unlike earlier models, which were restricted to either extractive or abstractive frameworks, LLMs bring unprecedented flexibility by seamlessly integrating both paradigms. With their in-context learning and few-shot learning capabilities, LLMs are able to handle complex summarization tasks with minimal supervision [15, 16]. This flexibility enables them to switch between extractive, abstractive and hybrid summarization techniques, without the need for retraining or significant architectural modifications. In addition to their generative flexibility, LLMs excel in producing coherent, semantically rich summaries by leveraging pre-trained knowledge from vast text corpora. These advancements have not only improved summarization quality but have also set new benchmarks for the field.
@@
Despite these advancements, several surveys on ATS remain focused on conventional extractive and abstractive methods [11, 12, 13, 14]. These conventional approaches, are often constrained by predefined generative paradigms and specific architectures. In contrast, this survey expands the scope by introducing a new category of LLM-based summarization techniques, focusing on the latest advancements in LLM-based summarization techniques, which offer more adaptive and flexible generative paradigms.
To ensure a comprehensive collection of research in LLM-based ATS, we developed an automatic retrieval algorithm that combines keyword-based searches with LLM-based prompt techniques to filter relevant papers. This algorithm is capable of capturing papers that explicitly mention “summarization” in their titles, as well as those that may not explicitly mention the term but are highly relevant to ATS. The proposed retrieval method can be adapted for use in other domains as well.
@@
In this study, our objective is to provide a comprehensive survey of ATS techniques, with a particular focus on exploring LLM-based methods. To the best of our knowledge, this work is among the first to review and analyze LLM-based ATS techniques in depth. We aim to make the following key contributions:
@@
    Up-to-date Survey on ATS: We provide a comprehensive review of the latest advancements in both conventional ATS approaches (extractive and abstractive summarization) and LLM-based methods that offer paradigm flexibility.
@@
    Reviewing LLM-Based Methods: We offer a detailed review of recent LLM-based summarization techniques, focusing on how in-context learning, prompt engineering, and few-shot learning have reshaped ATS.
  • 3.
    Proposed Automated Algorithm for ATS Paper Collection: We present an automated retrieval algorithm that combines keyword-based searches with LLM-based prompting to efficiently collect and organize research papers on ATS topics. This algorithm streamlines the paper collection process and can be adapted for use in other fields. The source code for the algorithm is available on GitHub [^1] https://github.com/JinHanLei/RetrievePapers.
The remainder of this paper is organized as follows: Section 2 provides an overview of the background of Automatic Text Summarization (ATS), including its categorization and prior surveys. Section 3 defines key concepts in ATS and outlines the paper collection methodology. Section 4 details the data acquisition methods used in the ATS pipeline. Summarization modeling approaches are discussed in Section 6. Section 7 describes evaluation metrics for summarization, and Section 8 highlights ATS-based applications. Finally, Section 9 explores future directions in ATS.
@@ 2.1 History of Automatic Text Summarization
Early Statistical Era (Pre-2000s). During this period, summarization methods were predominantly based on statistical text features, such as word counts or measures of term frequency. One of the foundational approaches was introduced by Hans Peter Luhn in 1958 [17], who utilized hand-crafted features like cue word lists, title words, and term frequencies to calculate a “significance” score for each word. Sentences containing the most significant words were then extracted to form an abstract, marking one of the earliest models of extractive summarization.
@@
Building upon Luhn’s methodology, later research in the era incorporated more advanced statistical models such as TF-IDF (Term Frequency-Inverse Document Frequency) [1], Latent Semantic Analysis (LSA) [2], and BM25 (Best Matching 25) [3]. These models aimed to compute sentence-to-sentence similarity scores, shifting from Luhn’s individual word frequency-based approach to a more sophisticated analysis of relationships between sentences. By constructing sentence relationship maps, key sentences could be identified and selected through path traversal methods, ensuring the overall quality of the generated summaries [18, 19, 20, 21, 22, 23, 24]. However, these methods were limited in capturing the deeper meanings or context of the text, and thus struggled to generate coherent summaries for more complex and semantically rich texts [25].
@@
In contrast to the purely statistical approaches, another line of research during this era focused on structure-based methods to create abstractive summaries. These methods aimed to identify the most important information from the text by using templates, rules, trees, and ontologies. For example, Gerani et al. (2014) [26] proposed converting texts into a discourse tree structure, selecting the most significant sentences, and refining them using hand-crafted templates to generate high-quality abstractive summaries. Similarly, ontology-based approaches treated summarization as the extraction of key entities and their relationships. Researchers like Li (2015) and others [27, 28, 29] adopted ontology methods to extract key phrases, which were then combined with other techniques to generate abstractive summaries. However, these methods were limited in capturing the deeper semantics of the input texts, heavily relied on hand-crafted features, and were thus constrained in their ability to generate accurate and coherent summaries [25].
@@
Machine Learning and Early Neural Networks Era (2000s-2010s). Machine learning-based approaches learn from data [30], alleviating the need for manually defined rules used in statistical methods. These methods can be supervised, unsupervised, or semi-supervised. In supervised learning, algorithms such as SVM (Support Vector Machines) [4, 31, 32], Naïve Bayes Classification [5, 33, 34], and CRF (Conditional Random Fields) [6, 35, 36] are trained on labeled data to classify sentences as “summary” or “non-summary.” On the other hand, unsupervised techniques use clustering algorithms such as K-Means [37, 38], Hierarchical Clustering [7, 39, 40], and DBSCAN (Density-Based Spatial Clustering of Applications with Noise) [8, 41]. Semi-supervised methods bridge the gap by requiring less labeled data to train an effective classifier. Nevertheless, these methods still rely heavily on feature engineering, and their performance can be highly dependent on the selection of features, limiting their generalizability and quality of summarization generation.
@@
Deep Learning and Transformer Era (2010s-2020s). With the success of deep learning techniques in Natural Language Processing (NLP), various neural networks achieved remarkable results in tasks such as machine translation, voice recognition, and dialogue systems by offering a deeper understanding of language semantics. One of the most important category of neural networks for summarization is Sequence-to-Sequence (Seq2Seq) models, which processed input text as sequential tokens and extracted semantic information through recurrent hidden neurons, and recursively generate sequential tokens as summaries. Examples include Recurrent Neural Networks (RNNs) [42, 43, 44], Long Short-Term Memory (LSTM) [45, 46, 47], and Gated Recurrent Units (GRU) [48, 49]. Despite their improvements, RNN-based models faced challenges in maintaining coherence over long sequences due to difficulty in capturing long-range dependencies, which limited their ability to produce high-quality summaries for longer documents.
@@
With the introduction of the Transformer neural network, the NLP landscape shifted from a “task-specific” paradigm to a “pre-trained” paradigm. Unlike previous models that required specific design adjustments for different NLP tasks, transformer-based language models are pre-trained on extensive text datasets and then fine-tuned for various objectives. On the architectural level, transformers feature sophisticated layers, such as multi-head attention, and consist of multiple neural layers, enabling them to capture deeper semantic meanings in texts. The summarization tasks are not exception, transformer-based neural models had became a “backbone” model at the current stage, such as BERT [50, 51, 52, 53, 54, 55], GPT [56, 57, 58], Sentence-BERT [59, 60], RoBERTa [61, 62, 63] and XLNET [64, 65, 66]. However, the major limitation of these pre-trained methods are 1) they require task-specific fine-tuning, which limits their scalability to training datasets and 2) the pretrained models generally lack flexibility in generation.
@@
In contrast to earlier approaches, which were constrained by a fixed summarization paradigm, either extractive or abstractive, LLMs offer unprecedented flexibility in their generative capabilities. Thanks to their robust text generation abilities, LLMs can seamlessly adapt between extractive and abstractive paradigms, or even blend both approaches within a single task. Moreover, due to their advanced language modeling, LLM-generated summaries tend to exhibit superior coherence, fluency, and overall writing quality compared to traditional methods.
LLM-based summarization methods can be broadly categorized into three key approaches:
@@
    Prompt Engineering-Based Methods: These methods involve designing effective prompts or templates to guide the LLMs in generating high-quality summaries without modifying the model’s internal parameters.
@@
    Fine-Tuning-Based Methods: In this approach, LLMs are fine-tuned on dedicated summarization datasets. By adjusting the internal parameters of the model, fine-tuning allows LLMs to specialize in summarization, further improving their accuracy and relevance for this task.
@@
    Knowledge Distillation-Based Methods: These approaches aim to extract knowledge from LLMs to create smaller, task-specific models. By distilling the summarization ability of LLMs into smaller models, it is possible to retain high performance while improving efficiency for specific summarization tasks.
The field of Automatic Text Summarization (ATS) has undergone a development path across several distinct eras, from early statistical approaches to the advent of Large Language Models (LLMs). Conventional ATS approaches are constrained by a fixed summarization paradigm, either extractive or abstractive. The emergence of LLMs represents a significant breakthrough, offering a flexible, paradigm-agnostic approach that surpasses previous limitations in both extractive and abstractive summarization. As the field continues to evolve, we anticipate that LLM-based methods will not only refine and enhance summarization capabilities but will also pave the way for more adaptive, scalable, and efficient solutions. Future research will likely focus on further improving the interpretability, scalability, and domain adaptability of these models, cementing their role as the foundation for next-generation ATS systems.
@@ 2.2 Categorization of Automatic Text Summarization
Automatic Text Summarization (ATS) systems are conventionally classified as “Extractive”, “Abstractive”, and “Hybrid” based on their generation paradigms. Extractive methods focus on selecting original sentences or phrases from the input text to create summaries, while abstractive approaches generate new sentences that convey the same meaning but differ from the original text. Hybrid methods combine both techniques. Conventional methods are inflexible, as they are specifically designed and trained to perform either extraction or abstractive tasks. In contrast, approaches based on Large Language Models (LLMs) unify summarization tasks within a generative framework, where both extractive and abstractive summarization can be achieved by prompting the LLM to directly generate the summarized text. Examples include having LLMs extract key information from the original text[67, 68], reading the original text and generating a summary[69, 70], and using hybrid methods[71, 72].
In this survey, we classify Automatic Text Summarization (ATS) systems into two major categories: Conventional ATS Methods and LLM-based ATS Methods. The first category encompasses conventional summarization approaches, including extractive, abstractive, and hybrid methods, which adhere strictly to their designated paradigms. In contrast, the second category includes the latest summarization techniques based on Large Language Models (LLMs), which demonstrate adaptability across various generation paradigms, enabling them to flexibly produce coherent summaries regardless of traditional constraints.
2.2.1 Conventional ATS Methods
  • 1.
    Extractive Methods selects important sentences or phrases directly from the original document to form a summary. It identifies key components in the text and uses them to create a coherent output [73]. Techniques include unsupervised methods, such as calculating and ranking the statistical importance of words and sentences, and supervised methods, which train machine learning or deep learning models to classify text as “summary” or “non-summary.”
  • 2.
    Abstractive Methods generates summaries by producing new sentences that may not appear in the original text. This can be achieved through structured methods, like tree or graph-based models, or generative approaches, such as sequence-to-sequence (seq2seq) architectures using RNNs [74] or Transformers [75]. Pre-trained models often enhance these methods, enabling more flexible and human-like summary generation.
  • 3.
    Hybrid Methods combines extractive and abstractive approaches. Typically, a hybrid ATS system uses an extractive model to identify key sentences, followed by an abstractive model to refine or rewrite the extracted content. This process can take two forms: a simpler “Extractive to Shallow Abstractive” approach or a more sophisticated “Extractive to Abstractive” approach with a specifically trained abstractive model [76].
2.2.2 LLM-based Methods

### v3

@@ Abstract
The exponential growth of textual content on the internet, alongside vast archives of news articles, scientific papers, legal documents, and other domains, has made Automatic Text Summarization (ATS) increasingly important. ATS aims to create concise and accurate summaries, significantly reducing the effort required to process large volumes of text. Originating in the 1950s, ATS has evolved through several technical shifts, moving from statistical models to machine learning and deep learning approaches, and more recently to pre-trained models. Previous surveys have focused on conventional ATS methods, which are often constrained by the predefined generative paradigms. However, the advent of Large Language Models (LLMs) has introduced a paradigm-flexible approach to summarization. With their superior generative capabilities, in-context learning, and few-shot learning abilities, LLMs have demonstrated remarkable improvements in coherence, fluency, and overall summarization quality. In this survey, we provide a comprehensive review of both conventional ATS approaches and the latest advancements in LLM-based methods. Our contributions include: (1) offering an up-to-date survey of ATS; (2) reviewing the latest LLM-based summarization methods.
@@ 1 Introduction
The exponential growth of the World Wide Web has led to an overwhelming surge in textual data across various domains, including news articles, websites, user reviews, blogs, and social media. Additionally, vast text archives exist in specialized fields such as books, scholarly papers, legal documents, and biomedical records. This rapid expansion of data has far outpaced individuals’ ability to search, read, and process all relevant information. To address this challenge, Automatic Text Summarization (ATS) has emerged, leveraging methods from Natural Language Processing (NLP) and Information Retrieval (IR) alongside advanced techniques such as Machine Learning (ML), Deep Learning (DL), and Large Language Models (LLMs). ATS enables users to grasp key ideas efficiently, significantly reducing the time and effort required for document reading and comprehension.
@@
Historically, ATS has undergone several distinct eras of development. In the pre‑2000s, early approaches were primarily statistical, leveraging surface‑level features such as word frequency and sentence position to generate summaries [1, 2, 3]. From the 2000s to the 2010s, machine learning and shallow neural network–based methods emerged, automating the feature extraction process but still heavily relying on domain‑specific, handcrafted features [4, 5, 6, 7, 8]. The introduction of deep learning in the 2010s—particularly encoder‑decoder architectures like RNNs [9] and LSTMs [10], enabled more nuanced, context‑aware summarization by modeling sequences of words and capturing richer semantic information. During this period, two main paradigms dominated: extractive summarization, which focused on identifying key sentences or phrases from the original text [11, 12], and abstractive summarization, which generated new sentences to paraphrase the main ideas [13, 14]. However, these approaches were constrained by their inflexible generative paradigms, which were determined by the specific model architecture and training data. Moreover, abstractive summarization often faced challenges in maintaining content coherence and ensuring factual accuracy.
@@
The advent of Large Language Models (LLMs), such as GPT‑3 and ChatGPT, represents a significant breakthrough in ATS by offering a paradigm‑flexible approach. Unlike earlier models restricted to extractive or abstractive frameworks, LLMs bring unprecedented flexibility by seamlessly integrating both paradigms. With their in‑context and few‑shot learning capabilities, LLMs can handle complex summarization tasks with minimal supervision [15, 16]. This flexibility enables them to switch between extractive, abstractive, and hybrid summarization techniques without retraining or significant architectural changes. In addition to this generative flexibility, LLMs excel at producing coherent, semantically rich summaries by leveraging pre‑trained knowledge from vast text corpora. These advancements have not only improved summarization quality but also set new benchmarks for the field.
@@
Despite these innovations, several surveys on ATS remain focused on conventional extractive and abstractive methods [11, 12, 13, 14]. These approaches are constrained by predefined generative paradigms and fixed architectures. In contrast, this survey introduces a new class of LLM‑based methods that offer more adaptive and flexible generative paradigms.
@@
In this study our objective is to provide a comprehensive survey of ATS techniques, with particular focus on LLM-based methods. We have developed an automated retrieval algorithm that combines keyword-based searches with LLM-based prompting to efficiently collect and organize research papers on ATS topics. This algorithm streamlines the paper collection process and can be adapted for use in other fields. The source code for the algorithm is available on GitHub. [^1] https://github.com/JinHanLei/RetrievePapers To the best of our knowledge, this work is among the first to review and analyze LLM-based ATS techniques in depth. We aim to make the following key contributions:
@@
    An Up‑to‑Date Survey of ATS: We provide a comprehensive review of the latest advancements in both conventional ATS approaches (extractive and abstractive summarization) and LLM-based methods that offer paradigm flexibility.
@@
    A Detailed Review of LLM‑Based Methods: We present an in‑depth analysis of recent LLM-based summarization techniques, focusing on how in‑context learning, prompt engineering, and few‑shot learning have reshaped ATS.
The remainder of this paper is organized as follows: Section 2 provides background on Automatic Text Summarization (ATS), including its categorization and prior surveys. Section 3 defines key concepts in ATS and outlines the paper collection methodology. Section 4 details the data acquisition methods used in the ATS pipeline. Conventional ATS approaches are discussed in Section 6. Section 7 explores the recent LLM-based summarization approaches. Section 8 describes evaluation metrics for summarization, and Section 9 highlights ATS-based applications. Finally, Section 10 outlines future directions in ATS.
@@ 2.1 History of Automatic Text Summarization
Early Statistical Era (Pre‑2000s). During this period, summarization methods were predominantly based on statistical text features such as word counts and term frequency measures. One of the foundational approaches was introduced by Hans Peter Luhn in 1958 [17]. He utilized hand‑crafted features like cue word lists, title words, and term frequencies to calculate a “significance” score for each word. Sentences containing the highest‑scoring words were then extracted to form an abstract, representing one of the earliest extractive summarization models.
@@
Building upon Luhn’s methodology, later research during this era incorporated more advanced statistical models such as TF‑IDF (Term Frequency‑Inverse Document Frequency) [1], Latent Semantic Analysis (LSA) [2], and BM25 (Best Matching 25) [3]. Researchers constructed sentence‑relationship maps to compute sentence similarity scores and then applied path traversal methods to select key sentences, ensuring the overall quality of generated summaries [18, 19, 20, 21, 22, 23, 24]. However, these approaches struggled to capture deeper context, resulting in summaries that often lacked coherence for semantically rich texts [25].
@@
In contrast to purely statistical methods, another line of research during this era focused on structure‑based approaches for abstractive summarization. These techniques used templates, rules, trees, and ontologies to identify and reformulate the most important information. For example, [26] converted texts into discourse trees, selected the most significant nodes, and applied hand‑crafted templates to generate abstractive summaries. Similarly, ontology‑based methods treated summarization as extracting key entities and relationships. Researchers such as Li [27] applied ontologies to extract essential phrases, which were then combined with template‑based techniques to produce coherent abstracts. These structure‑based approaches, however, depended heavily on hand‑crafted features and were limited in capturing deep semantics, constraining their ability to generate accurate and coherent summaries [25].
@@
Machine Learning and Early Neural Networks Era (2000s–2010s). Machine learning–based approaches learned from data [28], alleviating the need for manually defined rules used in statistical methods. These methods could be supervised, unsupervised, or semi‑supervised. In supervised learning, algorithms such as SVM (Support Vector Machines) [4, 29, 30], Naïve Bayes Classification [5, 31, 32], and CRF (Conditional Random Fields) [6, 33, 34] were trained on labeled data to classify sentences as “summary” or “non‑summary.” On the other hand, unsupervised techniques used clustering algorithms such as K‑Means [35, 36], Hierarchical Clustering [7, 37, 38], and DBSCAN (Density‑Based Spatial Clustering of Applications with Noise) [8, 39]. Semi‑supervised methods bridged the gap by requiring less labeled data to train an effective classifier. Nevertheless, these methods still relied heavily on feature engineering, and their performance could be highly dependent on feature selection, limiting their generalizability and the quality of generated summaries.
@@
Deep Learning and Transformer Era (2010s–2020s). With the success of deep learning techniques in Natural Language Processing (NLP), various neural networks achieved remarkable results in tasks such as machine translation, voice recognition, and dialogue systems by offering a deeper understanding of language semantics. One of the most important categories of neural networks for summarization is Sequence‑to‑Sequence (Seq2Seq) models, which processed input text as sequential tokens, extracted semantic information through recurrent hidden neurons, and recursively generated sequential tokens as summaries. Examples include Recurrent Neural Networks (RNNs) [40, 41, 42], Long Short‑Term Memory (LSTM) [43, 44, 45], and Gated Recurrent Units (GRU) [46, 47]. Despite their improvements, RNN‑based models faced challenges in maintaining coherence over long sequences due to difficulty in capturing long‑range dependencies, which limited their ability to produce high‑quality summaries for longer documents.
@@
With the introduction of the Transformer neural network, the NLP landscape shifted from a “task‑specific” paradigm to a “pre‑trained” paradigm. Unlike previous models that required specific design adjustments for different NLP tasks, transformer based language models are pre‑trained on extensive text datasets and then fine‑tuned for various objectives. At the architectural level, transformers feature sophisticated layers such as multi‑head attention and multiple neural layers, enabling them to capture deeper semantic meanings in texts. Summarization tasks are no exception: transformer‑based neural models have become the backbone for current methods, including BERT [48, 49, 50, 51, 52, 53], GPT [54, 55, 56], Sentence‑BERT [57, 58], RoBERTa [59, 60, 61], and XLNet [62, 63, 64]. However, the major limitations of these pre‑trained methods are: (1) they require task‑specific fine‑tuning, which limits scalability across datasets; and (2) they generally lack flexibility in generation.
@@
Unlike earlier approaches constrained to a single summarization paradigm—either extractive or abstractive, LLMs offer unprecedented flexibility in their generative capabilities. Thanks to their robust text generation abilities, LLMs can seamlessly adapt between extractive and abstractive paradigms or even blend both approaches within a single task. Moreover, LLM-generated summaries tend to exhibit superior coherence, fluency, and overall writing quality compared to traditional methods.
LLM-based summarization methods can be broadly categorized into four key approaches:
@@
    Prompt Engineering–Based Methods These methods involve designing effective prompts or templates to guide the LLMs in generating high-quality summaries without modifying the model’s internal parameters.
@@
    Retrieval-Augmented Generation–Based Methods These methods enhance the summarization process by incorporating externally retrieved documents to supplement the model’s internal knowledge.
@@
    Fine-Tuning–Based Methods In this approach, LLMs are fine-tuned on dedicated summarization datasets. By adjusting the model’s internal parameters, fine-tuning allows LLMs to specialize in summarization, further improving their accuracy and relevance for this task.
  4. 4.
    Knowledge Distillation–Based Methods These approaches extract knowledge from LLMs to create smaller, task-specific models. By distilling summarization capabilities into compact models, it is possible to retain high performance while improving efficiency for specific summarization tasks.
The field of Automatic Text Summarization (ATS) has followed a development path through several distinct eras, from early statistical approaches to the advent of Large Language Models (LLMs). Conventional ATS approaches are constrained by a fixed summarization paradigm, either extractive or abstractive. The emergence of LLMs represents a significant breakthrough, offering a flexible, paradigm-agnostic approach that surpasses previous limitations in both extractive and abstractive summarization. As the field continues to evolve, we anticipate that LLM-based methods will not only enhance summarization capabilities but also pave the way for more adaptive, scalable, and efficient solutions. Future research will likely focus on improving the interpretability, scalability, and domain adaptability of these models, cementing their role as the foundation for next-generation ATS systems.
@@ 2.2 Categorization of Automatic Text Summarization
Automatic Text Summarization (ATS) systems are conventionally classified as “Extractive”, “Abstractive” and “Hybrid” based on their generation paradigms. Extractive methods focus on selecting original sentences or phrases from the input text to create summaries, while abstractive approaches generate new sentences that convey the same meaning but differ from the original. Hybrid methods combine both techniques. Conventional methods are inflexible, as they are specifically designed and trained to perform either extraction or abstraction.
By contrast, LLM-based approaches unify summarization within a generative framework, enabling both extractive and abstractive outcomes through prompting. Examples include LLMs extracting key information from the original text [65, 66], reading the text and generating a summary [67, 68], and applying hybrid strategies [69, 70].
In this survey, we classify ATS methods into two major categories: Conventional Methods (extractive, abstractive and hybrid) and LLM-based Methods. The former adhere strictly to fixed paradigms, while the latter demonstrate adaptability across generation paradigms, flexibly producing coherent summaries without traditional constraints.
@@
2.2.1 Conventional ATS Methods
  • 1.
    Extractive Methods select important sentences or phrases directly from the original document to form a summary. It identifies key components in the text and uses them to create a coherent output [87]. Techniques include unsupervised methods, such as calculating and ranking the statistical importance of words and sentences, and supervised methods, which train machine learning or deep learning models to classify text as “summary” or “non-summary.”
  • 2.
    Abstractive Methods generate summaries by producing new sentences that may not appear in the original text. This can be achieved through structured methods, like tree or graph-based models, or generative approaches, such as sequence-to-sequence (seq2seq) architectures using RNNs [88] or Transformers [89]. Pre-trained models often enhance these methods, enabling more flexible and human-like summary generation.
  • 3.
    Hybrid Methods combine extractive and abstractive approaches. Typically, a hybrid ATS system uses an extractive model to identify key sentences, followed by an abstractive model to refine or rewrite the extracted content. This process can take two forms: a simpler “Extractive to Shallow Abstractive” approach or a more sophisticated “Extractive to Abstractive” approach with a specifically trained abstractive model [76].
2.2.2 LLM-based Methods